OpenStudy (risingtotop):

Your friend has $90 when he goes to the fair. He spends $5 to enter the fair and $15 on food. Rides at the fair cost $1.50 per ride. Which function can be used to determine how much money he has left over after x rides? (1 point) f(x) = 1.5x + 70 f(x) = -1.5x + 70 f(x) = -1.5x - 70 f(x) = -20x + $90

OpenStudy (brooke..help00):

Are you in FLVS?

OpenStudy (risingtotop):

?

OpenStudy (brooke..help00):

Ok, guess not anyways:) 5 + 15 = 20 (Just to get in and get food) 90-20= 70 (What you have left. ) So 1.5x (For each ride) f(x)= ? (With this info what do you think)

OpenStudy (risingtotop):

D

OpenStudy (risingtotop):

its A

OpenStudy (risingtotop):

@AloneS

alones (alones):

My gawd im confused..so your question was answered already?

OpenStudy (risingtotop):

no

OpenStudy (risingtotop):

she helped out but didnt tell if i was right

OpenStudy (risingtotop):

i think its A

alones (alones):

alright tho what about D?

OpenStudy (risingtotop):

its incorrect

OpenStudy (risingtotop):

hello?

alones (alones):

Alright lol so what do you think?

OpenStudy (risingtotop):

A

alones (alones):

Hrmm why A? Brooke got you trhue the steps Lol

alones (alones):

*through

OpenStudy (risingtotop):

yeah but am i correct?!?!!!!!!!!!!!!!!!!!!!!!!!!!!?!?!?!?!

alones (alones):

I'm not sure hwo you got A..But i got \(\ f(x) = -1.5x + 70\) at some point; Not sure if it's one of your choices tho ;3

OpenStudy (risingtotop):

wait how!??!? -1.5x

OpenStudy (risingtotop):

can you stop folling around honestly im trying to learn...

OpenStudy (risingtotop):

please :)

alones (alones):

Can you learn please (; Lol im not follin? what's that ...you mean fooling? LOL

alones (alones):

Hrmm so how did you get \(\ -5/x\) boi

OpenStudy (risingtotop):

😡😡😡😡😡😡😡😡v

alones (alones):

umm what's that^

OpenStudy (risingtotop):

you.....HAHHAHA

OpenStudy (risingtotop):

LOL

OpenStudy (risingtotop):

Get Rekt

OpenStudy (risingtotop):

😣

alones (alones):

You wanna laugh and see what im laughing at? http://prnt.sc/cw1l23

OpenStudy (risingtotop):

well they are so post to be emoji

alones (alones):

.-.

alones (alones):

I don't feel rekt

OpenStudy (risingtotop):

can you please please help me?

OpenStudy (risingtotop):

you're the best you're everything positive .... etc.... can you help <3

alones (alones):

Uhh i already said I got B Lol..i don't understand why you got A .-.

OpenStudy (risingtotop):

exactly explain

OpenStudy (risingtotop):

how did you get -1.5

OpenStudy (brooke..help00):

5 + 15 = 20 (Just to get in and get food) 90-20= 70 (What you have left. ) So 1.5x (For each ride) -1.5 (Subtract 1.5 for every ride) f(x) = -1.5x + 70

alones (alones):

EXACTLY!!!

OpenStudy (risingtotop):

O.O

OpenStudy (risingtotop):

I'm so confused

alones (alones):

Ugh I hope your eyes opened .-.

OpenStudy (risingtotop):

nope

alones (alones):

Oh my fluttering lord..what step you got confused on?

OpenStudy (risingtotop):

why negative

OpenStudy (risingtotop):

i understand why but why

OpenStudy (risingtotop):

OOOOOOOOOOOOOOO

OpenStudy (risingtotop):

I GET IT!!!!!!!!!!!!!!!!

alones (alones):

what's that "oooo" for?

OpenStudy (risingtotop):

IM SOOOOO STUPID

OpenStudy (brooke..help00):

Yay

alones (alones):

Oh Halleluja my child!

OpenStudy (risingtotop):

JK HAH

alones (alones):

Lmao you're serious or what?..-.

OpenStudy (risingtotop):

No i get the problem i mean giving you a medal lmao

alones (alones):

Oh wow LOL

OpenStudy (risingtotop):

haha you got scared

OpenStudy (risingtotop):

you probably thought "how much stupider can this kid get" but im like NOPE im not giving up

alones (alones):

It's fine..half of those users don't give medals ;-;

OpenStudy (risingtotop):

well too bad i gave you one

OpenStudy (risingtotop):

so deal with it

OpenStudy (risingtotop):

:)

likeabossssssss (likeabossssssss):

ok

alones (alones):

Alright ill try my best to deal with you LOL

alones (alones):

I wasted my time with this..glad your eyes opened at some point._.

OpenStudy (risingtotop):

A student is training for the swim team during July and August one summer. She swims one hour every day while averaging 800 meters each day. Select the graph that correctly shows the total meters swam during the time of her training program. (2 points)

alones (alones):

Oh lord is that a continue of your question?

likeabossssssss (likeabossssssss):

@RisingToTop plz but your q in difernt post ok

OpenStudy (risingtotop):

no its a differant question .-.

likeabossssssss (likeabossssssss):

thank u

alones (alones):

Oh alright make a new post hun.

alones (alones):

P.S dont' tag mee

likeabossssssss (likeabossssssss):

haha